2. Aqueous solutions of silver nitrate and potassium chloride can be mixed to precipitate silver


chloride, Calculate the mass of silver chloride that is formed in the reaction of 9.20 g silver


nitrate with 5.15 g potassium chloride.


1
Expert's answer
2022-03-08T07:59:01-0500

AgNO3+KCl->AgCl+KNO3

M(AgNO3)=108+14+16x3=170 g/mol

M(KCl)=39+35.5=74.5 g/mol

n(AgNO3)=9.2/170=0.0541 mol

n(KCl)=5.15/74.5=0.0691 mol

Limited component is AgNO3

M(AgCl)=108+35.5=143.5 g/mol

m(AgCl)=143.5x0.0541=7.76 g

Answer: 7.76 g of AgCl